OpenStudy (anonymous):

Find the fifth term of (2x + y)^10

OpenStudy (mathmate):

\((A+B)^n=nC0 A^0B^n+nC1 A^1B^{n-1}+...nCnA^nB^0\) So the 5th term is \(nC4A^4B^{n-4}\) where nCr="n choose r"=n!/(r!(n-r)!) For the given question, A=2x, B=y.